OpenStudy (anonymous):

factor 4x^2 - 4x + 1 = 0

OpenStudy (anonymous):

I know the outcome would be (2x-1)=0, but what is the method?

OpenStudy (yanasidlinskiy):

What do you mean by method??

OpenStudy (anonymous):

How exactly do you get it, how do you factor it.

OpenStudy (anonymous):

I completely forgot how to factor these types of problems.

OpenStudy (anonymous):

Complete the square

OpenStudy (anonymous):

Ah, alright. Do you think you can show me how to do it again?

OpenStudy (yanasidlinskiy):

This website might be a little helpful! http://www.mathsisfun.com/algebra/completing-square.html

OpenStudy (anonymous):

You need factors of 4*1 = 4 that sum to -4. What are they? and I'll help you with the next step.

OpenStudy (anonymous):

Well that would be just -2.

OpenStudy (anonymous):

Right, -2 * -2 = 4. since the factors are the same, you know you have a perfect square trinomial, which factors into the same factor, twice.

OpenStudy (anonymous):

Thefactors of 4x^2 are 2x and 2x; the factors of +1 are -1 and -1, so the trinomial factors into (2x-1)(2x-1) or (2x-1)^2

OpenStudy (anonymous):

Each factor produces the same root (solution) when set equal to zero, so you just have to solve 2x - 1 = 0 by adding 1 and dividing by 2.

OpenStudy (anonymous):

Perfect Square Factoring Formulas a² - 2ab + b² = (a - b)² a² + 2ab + b² = (a + b)²

OpenStudy (anonymous):

Ah alright. That makes more sense. I got confused for a second.

OpenStudy (anonymous):

In your case, a is 2x, and b is 1 in the top formula

OpenStudy (anonymous):

Alright thanks man.

OpenStudy (anonymous):

ur welcome